One integer is 9 more than another. Their product is 580. Find the integers?
Answer by TimothyLamb(4379)

y = x + 9
xy = 580
x(x + 9) = 580
xx + 9x - 580 = 0
---
the above quadratic equation is in standard form, with a=1, b=9 and c=-580
---
to solve the quadratic equation, by using the quadratic formula, copy and paste this:
1 9 -580
into this solver: https://sooeet.com/math/quadratic-equation-solver.php
---
the quadratic has two real roots at:
---
x = 20
x = -29
---
the negative root doesn't fit the problem statement, so use the positive root:
---
x = 20
y = x + 9
y = 20 + 9
y = 29
---
answer:
x = 20
y = 29
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php